A 2 mL solution containing 1.20 μCi/mL is injected into the blood stream of a patient. After dilution, a 1.00mL sample is withdrawn and found to have an activity of 2.4×10−4 μCi . Calculate the total blood volume in Liters
Answer #1

Rsample = Rtracer * (Wsample / (Wsystem + Wtracer))

soo

2.4*10 ^-4 = 1.20 * (1/(Wsystem + 2))

Solve for Wsystem

2.4*10 ^-4 = 1.20 / (Wsystem + 2)

Wsystem = 1.20 / (2.4*10 ^-4) - 2

Wsystem = 4998 mL